please help
for the following equation state the value of the discriminant and then describe the nature of the solutions
13x^2-4x-10=0
thanks

Answer by stanbon(75887)

for the following equation state the value of the discriminant and then describe the nature of the solutions
13x^2-4x-10=0
----
b^2-4ac = 4^2-4*13*-10 >0 so
there are 2 unequal Real Number solutions.
=============================================
Cheers,
Stan H.
============